# Refereed Learning

### v1

@@ 4.1 Refereed learning for zero-one loss
List of Protocols 5 refereed learning for zero-one loss
First, we argue that Protocol 5 satisfies the soundness condition of Definition 4.1. Assume without loss of generality that $\mathcal{L}_{\mathcal{D}}(h_{1},f\mid\ell_{\mathsf{zo}})>(1+\varepsilon)\cdot\mathcal{L}_{\mathcal{D}}(h_{0},f\mid\ell_{\mathsf{zo}})$. Since $\ell_{\mathsf{zo}}$ is the zero-one metric, this is equivalent to the assumption that $\operatorname*{{Pr}}_{\mathcal{D}}{\left[{h_{1}(x)\neq f(x)}\right]}>(1+\varepsilon)\cdot\operatorname*{{Pr}}_{\mathcal{D}}{\left[{h_{0}(x)\neq f(x)}\right]}$. To prove that the verifier in Protocol 5 outputs the correct bit, we prove 4.3, which roughly states that for $x\sim{\mathcal{D}}|_{S}$, we have $h_{1}(x)\neq f(x)$ with probability at least $\frac{1}{2}+\frac{\varepsilon}{1+\varepsilon}$.
@@
To complete the proof of Theorem 4.2, observe that by Lemma 3.2, the verifier correctly obtains $p_{S}={\mathcal{D}}(S)$. Since ${\mathcal{D}}|_{S}(x)=\frac{{\mathcal{D}}(x)\cdot\mathds{1}{\left[{x\in S}\right]}}{{\mathcal{D}}(S)}$, the verifier can provide query access to $Q_{{\mathcal{D}}|_{S}}$, the probability mass function of ${\mathcal{D}}|_{S}$. Thus, by Lemma 3.1, we have $d_{\mathrm{TV}}{\big({\widehat{{\mathcal{D}}},{\mathcal{D}}|_{S}}\big)}\leq\delta$ and therefore, by 4.3 and the definition of $S$, we have
@@
To see why the verifier outputs 0 with probability at least $1-\beta$, let $\widehat{p}=\frac{1}{m}\sum_{i\in[m]}\mathds{1}{\left[{h_{0}(x_{i})\neq f(x_{i})}\right]}$. Since $\mathop{{\mathds{E}}\/}[\widehat{p}]=\operatorname*{{Pr}}[h_{0}(x)\neq f(x)]$, Hoeffding’s inequality and our setting of $m$ in Protocol 5 implies
@@
It follows that $\widehat{p}<\frac{1}{2}-\frac{\varepsilon}{2(2+\varepsilon)}+2\delta\leq\frac{1}{2}$ with probability at least $1-\beta$, and therefore $\mathcal{V}$ will output $\rho=0$ with probability at least $1-\beta$. The runtime and communication complexity follow by inspection of Protocol 5, and from the guarantees of Lemmas 3.1 and 3.2. ∎
@@ 4.2 Refereed learning for metric loss functions
We define the protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} in Protocol 6.
@@
List of Protocols 6 refereed learning for metric loss
@@
is less than $\frac{1}{2}-\varepsilon$ whenever $\mathcal{L}_{\mathcal{D}}(h_{1-b},f\mid\ell)>(3+\varepsilon)\cdot\mathcal{L}_{\mathcal{D}}(h_{b},f\mid\ell)$. Then, we will argue that the estimate $\widehat{R}_{b}$ computed by $\mathcal{V}$ in Protocol 6 is concentrated around $\mathop{{\mathds{E}}\/}{\left[{\widehat{R}_{b}}\right]}$, and that $R_{b}$ and $\mathop{{\mathds{E}}\/}{\left[{\widehat{R}_{b}}\right]}$ are close together. Combining the above with the fact that $R_{0}+R_{1}=1$ suffices to complete the proof.
@@
Now, since $r_{b}(x)\in[0,1]$ and each $x_{i}$ in Protocol 6 is sampled from $\widehat{{\mathcal{D}}}$ independently, Hoeffding’s inequality implies that
@@
Next, we bound the distance between $R_{b}$ and $\mathop{{\mathds{E}}\/}{\left[{\widehat{R}_{b}}\right]}$. Recall that $R_{b}=\mathop{{\mathds{E}}\/}_{x\sim{\mathcal{D}}^{h_{0},h_{1}}_{\ell}}{\left[{r_{b}(x)}\right]}$, and that $\mathop{{\mathds{E}}\/}{\left[{\widehat{R}_{b}}\right]}=\mathop{{\mathds{E}}\/}_{x\sim\widehat{{\mathcal{D}}}}{\left[{r_{b}(x)}\right]}$. Since $r_{b}(x)\in[0,1]$ and $d_{\mathrm{TV}}{\big({\widehat{{\mathcal{D}}},{\mathcal{D}}^{h_{0},h_{1}}_{\ell}}\big)}\leq\delta$, we have
@@
Combining (4) and the above bound yields ${\left|{\widehat{R}_{b}-R_{b}}\right|}\leq 2\delta$ for each $b\in\{0,1\}$ with probability at least $1-\beta$. To complete the proof, suppose $\mathcal{L}_{\mathcal{D}}{\left({h_{1-s},f}\right)}>(3+\varepsilon)\mathcal{L}_{\mathcal{D}}{\left({h_{s},f}\right)}$ for some $s\in\{0,1\}$. By 4.6 we have $R_{s}<\frac{1}{2}-\frac{\varepsilon}{2(2+\varepsilon)}$, and since $R_{0}+R_{1}=1$, this implies that $R_{1-s}>\frac{1}{2}+\frac{\varepsilon}{2(2+\varepsilon)}$. If $|\widehat{R}_{b}-R_{b}|\leq 2\delta$ for each $b\in\{0,1\}$ then by our choice of $\delta$ we have $\widehat{R}_{s}<\frac{1}{2}<\widehat{R}_{1-s}$. Since $\rho=\arg\min_{b\in\{0,1\}}\widehat{R}_{b}$, we see that the verifier outputs $\rho=s$ with probability at least $1-\beta$. The runtime, communication complexity, and query complexity guarantees follow from Lemmas 3.2 and 3.1, and by inspection of Protocol 6. ∎
@@ 5.1 Verification with labeled samples
Now, for each $b\in\{0,1\}$ and $f\leftarrow h_{b}$, define the view of the verifier $\textsf{view}{\left({\mathcal{V}^{{\mathcal{D}}_{f}}}\right)}_{b}$ as the distribution over the $m$ samples $(x,f(x))$ drawn from ${\mathcal{D}}_{f}$, and the transcripts $T_{b}$ and $T_{1-b}$ between $\mathcal{V}$ and ${\mathcal{P}}^{f}_{b}$, and between $\mathcal{V}$ and ${\mathcal{P}}^{*}_{1-b}$. Let $E$ be the event that one of the $m$ samples drawn by the verifier is $(0^{d},f(0^{d}))$. Since ${\mathcal{P}}^{*}_{1-b}$ executes ${\mathcal{P}}^{h_{1-b}}_{1-b}$ and the honest prover executes ${\mathcal{P}}^{f}_{b}={\mathcal{P}}^{h_{b}}_{b}$, the distribution of ($T_{0},T_{1})$ is independent of $b$ (the verifier always interacts with ${\mathcal{P}}^{h_{0}}_{0}$ and ${\mathcal{P}}^{h_{1}}_{1}$), and thus $\textsf{view}{\left({\mathcal{V}^{{\mathcal{D}}_{f}}\mid\overline{E}}\right)}_{0}=\textsf{view}{\left({\mathcal{V}^{{\mathcal{D}}_{f}}\mid\overline{E}}\right)}_{1}$. Next, we utilize the following fact from [RS06].
Fact 5.2 (Claim 4 [RS06]).
@@ 6.1 Efficient refereed learning for juntas
Proof.
@@
The soundness of the verifier follows by the same argument as in the proof of Theorem 4.2. The runtime and query complexity of the verifier, and the communication complexity of the protocol follows by Lemmas 3.2 and 3.5. ∎
@@ 6.2 When $\mathcal{D}$} and $\ell$ are arbitrarily precise
We now explain how our protocols can be implemented for distributions $\mathcal{D}$} and metrics $\ell$ that are not $\lambda$-precise (Definition 2.4); however, in order to bound communication complexity, we still require $\ell$ to be $M$-bounded—that is, $\ell(y,y^{\prime})\leq M$ for all $y,y^{\prime}\in\mathcal{Y}$. Specifically, we analyze the cost of rounding $Q_{\mathcal{D}}$ and $\ell$ so that they are $\lambda$-precise. Combined with the protocols in Section 4, this allows us to design protocols with additive error $\eta$ for arbitrarily precise distributions and loss functions that only incur communication cost $\log\frac{1}{\eta}\mathop{\mathrm{poly}}d$ (there is no cost in query complexity).
Recall that by 3.7, for all distributions $\mathcal{D}$} over $\{0,1\}^{d}$, the distribution ${\mathcal{D}}_{\lambda}$ (defined by ${\mathcal{D}}_{\lambda}(x)=\frac{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}{\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}$, where $\lfloor{y}\rfloor_{\lambda}$ denotes $2^{-\lambda}\cdot\lfloor{2^{\lambda}\cdot y}\rfloor$ for all $y\in\mathbb{R}$—that is, $\lfloor{y}\rfloor_{\lambda}$ denotes the nearest multiple of $2^{-\lambda}$ that is at most $y$) is $\lambda$-precise and satisfies $d_{\mathrm{TV}}({\mathcal{D}},{\mathcal{D}}_{\lambda})\leq 2^{d+1-\lambda}$. Now, for metric $\ell$, let $\ell_{\lambda}(y,y^{\prime})$ denote $\lfloor{\ell(y,y^{\prime})}\rfloor_{\lambda}$. Then, by definition of $\lfloor{\cdot}\rfloor_{\lambda}$ we have ${\left|{\ell_{\lambda}(y,y^{\prime})-\ell(y,y^{\prime})}\right|}\leq 2^{-\lambda}$. At a high level, the goal of this section is to explain how one can execute the protocols of Theorems 4.2 and 4.4 using ${\mathcal{D}}_{\lambda}$ and $\ell_{\lambda}$, given an $M$-bounded metric $\ell$ and query access to $Q_{\mathcal{D}}$, at the cost of a small additive error term $\eta$ and factor of $\log\frac{1}{\eta}$ in communication complexity and runtime.
In Proposition 6.2 we assume that for any metric $\ell$, computing $\lfloor{\ell(y,y^{\prime})}\rfloor_{\lambda}$ takes unit time.
@@
It remains to show that if $\mathcal{L}_{\mathcal{D}}(h_{1-b},f\mid\ell)>\alpha\mathcal{L}_{\mathcal{D}}(h_{b},f\mid\ell)+\eta$ for some $b\in\{0,1\}$, then $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} outputs $b$ with probability at least $1-\beta$. We will argue that the loss functions $\mathcal{L}_{\mathcal{D}}(f,h\mid\ell)$ and $\mathcal{L}_{{\mathcal{D}}_{\lambda}}(f,h\mid\ell_{\lambda})$ are close. Let $a(x)=\ell(f(x),h(x))-\ell_{\lambda}(f(x),h(x))$ and let $b(x)={\mathcal{D}}(x)-{\mathcal{D}}_{\lambda}(x)$ for all $x\in\{0,1\}^{d}$. Then by the definition of $\lfloor{\cdot}\rfloor_{\lambda}$ and 3.7, we have $|a(x)|\leq 2^{-\lambda}$ and $|b(x)|\leq 2^{d+1-\lambda}$, and thus,

### v2

@@ 4.1 Refereed learning for zero-one loss
We define the protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} in Protocol 4.
@@
List of Protocols 4 refereed learning for zero-one loss
First, we argue that Protocol 4 satisfies the soundness condition of Definition 4.1. Assume without loss of generality that $\mathcal{L}_{\mathcal{D}}(h_{1},f\mid\ell_{\mathsf{zo}})>(1+\varepsilon)\cdot\mathcal{L}_{\mathcal{D}}(h_{0},f\mid\ell_{\mathsf{zo}})$. Since $\ell_{\mathsf{zo}}$ is the zero-one metric, this is equivalent to the assumption that $\operatorname*{{Pr}}_{\mathcal{D}}{\left[{h_{1}(x)\neq f(x)}\right]}>(1+\varepsilon)\cdot\operatorname*{{Pr}}_{\mathcal{D}}{\left[{h_{0}(x)\neq f(x)}\right]}$. To prove that the verifier in Protocol 4 outputs the correct bit, we prove 4.3, which roughly states that for $x\sim{\mathcal{D}}|_{S}$, we have $h_{1}(x)\neq f(x)$ with probability at least $\frac{1}{2}+\frac{\varepsilon}{1+\varepsilon}$.
@@
To complete the proof of Theorem 4.2, observe that by Lemma 3.2, the verifier correctly obtains $p_{S}={\mathcal{D}}(S)$. Since ${\mathcal{D}}|_{S}(x)=\frac{{\mathcal{D}}(x)\cdot\mathds{1}{\left[{x\in S}\right]}}{{\mathcal{D}}(S)}$, the verifier has query access to $Q_{{\mathcal{D}}|_{S}}$, the probability mass function of ${\mathcal{D}}|_{S}$. Thus, by Lemma 3.1, we have $d_{\mathrm{TV}}{\big({\widehat{{\mathcal{D}}},{\mathcal{D}}|_{S}}\big)}\leq\delta$ and therefore, by 4.3 and the definition of $S$, we have
@@
To see why the verifier outputs 0 with probability at least $1-\beta$, let $\widehat{p}=\frac{1}{m}\sum_{i\in[m]}\mathds{1}{\left[{h_{0}(x_{i})\neq f(x_{i})}\right]}$. Since $\mathop{{\mathds{E}}\/}[\widehat{p}]=\operatorname*{{Pr}}[h_{0}(x)\neq f(x)]$, Hoeffding’s inequality and our setting of $m$ in Protocol 4 implies
@@
It follows that $\widehat{p}<\frac{1}{2}-\frac{\varepsilon}{2(2+\varepsilon)}+2\delta\leq\frac{1}{2}$ with probability at least $1-\beta$, and therefore $\mathcal{V}$ will output $\rho=0$ with probability at least $1-\beta$. The runtime and communication complexity follow by inspection of Protocol 4, and from the guarantees of Lemmas 3.1 and 3.2. ∎
@@ 4.2 Refereed learning for metric loss functions
List of Protocols 5 refereed learning for metric loss
@@
is less than $\frac{1}{2}-\varepsilon$ whenever $\mathcal{L}_{\mathcal{D}}(h_{1-b},f\mid\ell)>(3+\varepsilon)\cdot\mathcal{L}_{\mathcal{D}}(h_{b},f\mid\ell)$. Then, we will argue that the estimate $\widehat{R}_{b}$ computed by $\mathcal{V}$ in Protocol 5 is concentrated around $\mathop{{\mathds{E}}\/}{\left[{\widehat{R}_{b}}\right]}$, and that $R_{b}$ and $\mathop{{\mathds{E}}\/}{\left[{\widehat{R}_{b}}\right]}$ are close together. Combining the above with the fact that $R_{0}+R_{1}=1$ suffices to complete the proof.
@@
Now, since $r_{b}(x)\in[0,1]$ and each $x_{i}$ in Protocol 5 is sampled from $\widehat{{\mathcal{D}}}$ independently, Hoeffding’s inequality implies
@@
Next, we bound the distance between $R_{b}$ and $\mathop{{\mathds{E}}\/}{\left[{\widehat{R}_{b}}\right]}$. Recall $R_{b}=\mathop{{\mathds{E}}\/}_{x\sim{\mathcal{D}}^{h_{0},h_{1}}_{\ell}}{\left[{r_{b}(x)}\right]}$, and $\mathop{{\mathds{E}}\/}{\left[{\widehat{R}_{b}}\right]}=\mathop{{\mathds{E}}\/}_{x\sim\widehat{{\mathcal{D}}}}{\left[{r_{b}(x)}\right]}$. Since $r_{b}(x)\in[0,1]$ and $d_{\mathrm{TV}}{\big({\widehat{{\mathcal{D}}},{\mathcal{D}}^{h_{0},h_{1}}_{\ell}}\big)}\leq\delta$, we have
@@
Combining (4) and the above bound yields ${\left|{\widehat{R}_{b}-R_{b}}\right|}\leq 2\delta$ for each $b\in\{0,1\}$ with probability at least $1-\beta$. To complete the proof, suppose $\mathcal{L}_{\mathcal{D}}{\left({h_{1-s},f}\right)}>(3+\varepsilon)\mathcal{L}_{\mathcal{D}}{\left({h_{s},f}\right)}$ for some $s\in\{0,1\}$. By 4.6 we have $R_{s}<\frac{1}{2}-\frac{\varepsilon}{2(2+\varepsilon)}$, and since $R_{0}+R_{1}=1$, this implies that $R_{1-s}>\frac{1}{2}+\frac{\varepsilon}{2(2+\varepsilon)}$. If $|\widehat{R}_{b}-R_{b}|\leq 2\delta$ for each $b\in\{0,1\}$ then by our choice of $\delta$ we have $\widehat{R}_{s}<\frac{1}{2}<\widehat{R}_{1-s}$. Since $\rho=\arg\min_{b\in\{0,1\}}\widehat{R}_{b}$, we see that the verifier outputs $\rho=s$ with probability at least $1-\beta$. The runtime, communication complexity, and query complexity guarantees follow from Lemmas 3.2 and 3.1, and by inspection of Protocol 5. ∎
@@ 5.1 Verification with labeled samples
Now, for each $b\in\{0,1\}$ and $f\leftarrow h_{b}$, define the view of the verifier $\textsf{view}{\left({\mathcal{V}^{{\mathcal{D}}_{f}}}\right)}_{b}$ as the distribution over the $m$ samples $(x,f(x))$ drawn from ${\mathcal{D}}_{f}$, and the transcripts $T_{b}$ and $T_{1-b}$ between $\mathcal{V}$ and ${\mathcal{P}}^{f}_{b}$, and between $\mathcal{V}$ and ${\mathcal{P}}^{*}_{1-b}$. Let $E$ be the event that one of the $m$ samples drawn by the verifier is $(0^{d},f(0^{d}))$. Since ${\mathcal{P}}^{*}_{1-b}$ executes ${\mathcal{P}}^{h_{1-b}}_{1-b}$ and the honest prover executes ${\mathcal{P}}^{f}_{b}={\mathcal{P}}^{h_{b}}_{b}$, the distribution of ($T_{0},T_{1})$ is independent of $b$ (the verifier always interacts with ${\mathcal{P}}^{h_{0}}_{0}$ and ${\mathcal{P}}^{h_{1}}_{1}$), and thus $\textsf{view}{\left({\mathcal{V}^{{\mathcal{D}}_{f}}\mid\overline{E}}\right)}_{0}=\textsf{view}{\left({\mathcal{V}^{{\mathcal{D}}_{f}}\mid\overline{E}}\right)}_{1}$. Next, we utilize the following fact from [RaskhodnikovaS06].
Fact 5.2 (Claim 4 [RaskhodnikovaS06]).
@@ 6.1 Efficient refereed learning for juntas
To prove this statement, we use the following theorem, which can be viewed as an efficient, special case of certifiable sample (Lemma 3.1) for sampling uniformly from some set $S\subseteq\{0,1\}^{d}$.
Claim 6.2 (Certifiable index).
  1. 1.
  2. 2.
Proof of 6.2.
List of Protocols 6 certifiable index
Without loss of generality assume ${\mathcal{P}}_{0}$ is honest. Then $\widehat{x}=S_{i}$ and hence |A|=i1|A|=i-1. By Lemma 3.2, protocol ${\left[{{\mathcal{P}}_{0},{\mathcal{P}}^{*}_{1},\mathcal{V}}\right]}_{0}$ outputs accept. Next, consider ${\left[{{\mathcal{P}}_{0},{\mathcal{P}}^{*}_{1},\mathcal{V}}\right]}_{b}$. If ${\mathcal{P}}^{*}_{1}$ sends $\widehat{x}\neq S_{i}$, then, if $\widehat{x}\not\in S$, the protocol rejects in Step 2. On the other hand, if $\widehat{x}\in S$ then $|A|=|{\left\{{x\in S\mid x\prec\widehat{x}}\right\}}\neq i-1$. By Lemma 3.2, the sum $\sum_{x\in\{0,1\}^{d}}t(x)=|A|\neq i-1$, and hence the protocol outputs reject. To complete the proof, define the protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} as follows: for each $b\in\{0,1\}$ run protocols ${\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]}_{b}$, and return the $\widehat{x}$ from the first round of an accepting execution. By the above argument, the protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} always outputs $S_{i}$. The communication complexity and runtime guarantees follow by inspection of Protocol 6 and from the guarantees of Lemma 3.2. ∎
Proof of Proposition 6.1.
@@
The soundness of the verifier follows by the same argument as in the proof of Theorem 4.2. The runtime and query complexity of the verifier, and the communication complexity of the protocol follows by Lemmas 3.2 and 6.2. ∎
@@ 6.2 When $\mathcal{D}$} and $\ell$ are arbitrarily precise
We now explain how our protocols can be implemented for distributions $\mathcal{D}$} and metrics $\ell$ that are not $\lambda$-precise; however, in order to bound communication complexity, we still require $\ell$ to be $M$-bounded—that is, $\ell(y,y^{\prime})\leq M$ for all $y,y^{\prime}\in\mathcal{Y}$. Specifically, we analyze the cost of rounding $Q_{\mathcal{D}}$ and $\ell$ so that they are $\lambda$-precise. Combined with the protocols in Section 4, this allows us to design protocols with additive error $\eta$ for arbitrarily precise distributions and loss functions that only incur communication cost $\log\frac{1}{\eta}\mathop{\mathrm{poly}}d$ (there is no cost in query complexity).
Recall that by 3.3, for all distributions $\mathcal{D}$} over $\{0,1\}^{d}$, the distribution ${\mathcal{D}}_{\lambda}$ (defined by ${\mathcal{D}}_{\lambda}(x)=\frac{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}{\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}$, where $\lfloor{y}\rfloor_{\lambda}$ denotes $2^{-\lambda}\cdot\lfloor{2^{\lambda}\cdot y}\rfloor$ for all $y\in\mathbb{R}$) satisfies $d_{\mathrm{TV}}({\mathcal{D}},{\mathcal{D}}_{\lambda})\leq 2^{d+1-\lambda}$. Now, for metric $\ell$, let $\ell_{\lambda}(y,y^{\prime})$ denote $\lfloor{\ell(y,y^{\prime})}\rfloor_{\lambda}$. Then, by definition of $\lfloor{\cdot}\rfloor_{\lambda}$ we have ${\left|{\ell_{\lambda}(y,y^{\prime})-\ell(y,y^{\prime})}\right|}\leq 2^{-\lambda}$. At a high level, the goal of this section is to explain how one can execute the protocols of Theorems 4.2 and 4.4 using ${\mathcal{D}}_{\lambda}$ and $\ell_{\lambda}$, given an $M$-bounded metric $\ell$ and query access to $Q_{\mathcal{D}}$, at the cost of a small additive error term $\eta$ and factor of $\log\frac{1}{\eta}$ in communication complexity and runtime.
In Proposition 6.3 we assume that, for any metric $\ell$, computing $\lfloor{\ell(y,y^{\prime})}\rfloor_{\lambda}$ takes unit time.
@@
It remains to show that if $\mathcal{L}_{\mathcal{D}}(h_{1-b},f\mid\ell)>\alpha\mathcal{L}_{\mathcal{D}}(h_{b},f\mid\ell)+\eta$ for some $b\in\{0,1\}$, then $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} outputs $b$ with probability at least $1-\beta$. We will argue that the loss functions $\mathcal{L}_{\mathcal{D}}(f,h\mid\ell)$ and $\mathcal{L}_{{\mathcal{D}}_{\lambda}}(f,h\mid\ell_{\lambda})$ are close. Let $a(x)=\ell(f(x),h(x))-\ell_{\lambda}(f(x),h(x))$ and let $b(x)={\mathcal{D}}(x)-{\mathcal{D}}_{\lambda}(x)$ for all $x\in\{0,1\}^{d}$. Then by the definition of $\lfloor{\cdot}\rfloor_{\lambda}$ and 3.3, we have $|a(x)|\leq 2^{-\lambda}$ and $|b(x)|\leq 2^{d+1-\lambda}$, and thus,